Sarah has a deck of 52 cards and ramdomly selects 2. What is the probability that she will pick the same cards?
Answer by stanbon(75887)

Sarah has a deck of 52 cards and ramdomly selects 2. What is the probability that she will pick the same cards?
---
# of ways to succeed:
1st: pick a card type: 13 ways
2nd: pick 2 of that card: 4C2 = 6 ways
-----
# of random results:
Pick 2 cards from the deck: 52C2
=====
P(2 of the same card type) = (13*6)/52C2 = 78/1326 = 1/17
=================================
Cheers,
Stan H.
================